5. The economy below would be characterized as a

domestic output AE, closed economy government
200 230 20
250 270 20
300 310 20
350 350 20
400 390 20
450 430 20
500 470 20

A) private closed economy

B) private open economy

C) closed mixed economy

D) open mixed economy

6. If all forms of spending below are integrated into the economy, equilibrium GDP is

domestic output AE, closed economy government
200 230 20
250 270 20
300 310 20
350 350 20
400 390 20
450 430 20
500 470 20

A) 300

B) 350

C)400

D)450

7. If the below economy was closed to international trade, the equilibrium GDP is

domestic output AE, closed economy exports imports
200 230 30 20
250 270 30 20
300 310 30 20
350 350 30 20
400 390 30 20
450 430 30 20
500 470 30 20

A) $300

B) $350

C) $400

D) $450

8. Refer to the below table. The spending multiplier in this economy would be

domestic output AE, closed economy exports imports
200 230 30 20
250 270 30 20
300 310 30 20
350 350 30 20
400 390 30 20
450 430 30 20
500 470 30 20

A) 2

B) 3

C) 4

D) 5

9. Refer to the below table. For the open economy, the equilibrium GDP is

domestic output AE, closed economy exports imports
200 230 30 20
250 270 30 20
300 310 30 20
350 350 30 20
400 390 30 20
450 430 30 20
500 470 30 20

A) $300

B) $350

C) $400

D) $450

10. If net exports decline from zero to some negative amount, the aggregate expenditures schedule would

A) shift upward

B) shift downward
C) not move, as net exports do not affect aggregate expenditures
D) become steeper

11. If the spending multiplier in an economy is 5, a $20 billion decrease in taxes will

A) increase gdp by $80 billion

B) increase gdp by $100 billion
C) decrease gdp by $80 billion
D) decrease gdp by $100 billion

12. The tax multiplier for the below economy would be

domestic output AE, closed economy government
200 230 20
250 270 20
300 310 20
350 350 20
400 390 20
450 430 20
500 470 20

A) -2
B) -3

C) -4
D) -5

13. If the economy is operating at a point above the equilibrium point, then

A) AE is greater than GDP

B) GDP is greater than AE
C) The economy is at a sustainable full employment level
D) Business inventories have been depleted

14. Which of the following is NOT true in the Keynesian aggregate expenditures model?

A) Price level is held constant

B) There is a direct relationship between aggregate expenditures and GDP
C) The wealth effect has no bearing on the aggregate expenditures model

D) All changes in spending are autonomous and do not change the MPC

15. Other things equal, of a change in the tastes of American consumers causes them to purchase more foreign goods at each level of US GDP

A) Unemployment will decrease domestically

B) US GDP will fall
C) Inflation will occur domestically
D) US real GDP will rise
Answer #1

(5) (C)

Since AE values are for closed economy, this is a closed economy. Since value of government spending is given, this is a mixed economy.

(6) (D)

In equilibrium, Domestic output = AE for closed economy + Government spending, which is satisfied when

450 = 430 + 20

(7) (B)

In closed-economy equilibrium, Domestic output = AE for closed economy, which is satisfied when

350 = 350

(8) (D)

Marginal propensity to consume (MPC) = Change in Domestic output / Change in AE = (270 - 230) / (250 - 200) = 40/50

= 0.8

Multiplier = 1 / (1 - MPC) = 1 / (1 - 0.8) = 1/0.2 = 5

NOTE: As per Answering Policy, 1st 4 questions are answered.